OpenStudy (anonymous):

What is the equation of a parabola with a vertex at (0, 0) and a focus at (0, 3).

OpenStudy (anonymous):

standard form: x^2=4ay here a = 3 =>x^2=12y

OpenStudy (anonymous):

I don't want the standard form I need the conic one :/ like this: x = -1/8 y^2

OpenStudy (anonymous):

this is the standard form of parabola!

OpenStudy (anonymous):

is there any other forms?

OpenStudy (anonymous):

nope!